OpenStudy (anonymous):

If f(x)=x^3, then the inverse function f^-1(x)=?

OpenStudy (anonymous):

y=x^3 for inverse put y instead of x and x instead of y x=y^3 hence y= cube root of x answer f-1(x)=y=cube root of x

OpenStudy (anonymous):

im such a procrastinator, i had all spring break to do this homework and im just starting it NOW. thanks so much.